OpenStudy (anonymous):

Simplify the following expression, and rewrite it in an equivalent form with positive exponents. 6x^2y^3 24xy^2

OpenStudy (anonymous):

\[\LARGE \frac{6x^2y^3}{24xy^2}\] is it fraction ? :)

OpenStudy (anonymous):

how you write fractions kreshnik?

OpenStudy (anonymous):

\LARGE \frac { 6x^2 y^3 } { 24 x y^2 } copy paste this in Equation button... and you'll see :)

OpenStudy (turingtest):

@myko there are two ways: {x\over y} \frac{x}{y}

OpenStudy (anonymous):

that's right :)

OpenStudy (anonymous):

Dear asker, while you didn't confirm my question, if it's fraction or not... I'll write it like fraction... \[\LARGE \frac{6x^2y^3}{24xy^2}=\frac{x y}{4}=\frac14 \cdot xy\]